the perimeter of a rectangle is 36 m. if the length is three times the width, what is the width?

Answer by checkley79(3341)

P=2L+2W
L=3W
36=2*3W+2W
36=6W+2W
36=8W
W=36/8
W=4.5 ANS. FOR THE WIDTH.
L=3*2.5=7.5
PROOF:
36=2*3*4.5+2*4.5
36=6*4.5+9
36=27+9
36=36